OpenStudy (anonymous):

For the simple harmonic motion equation d=8 sin(4 pi t) , what is the period? If necessary, use the slash (/) to represent a fraction.

zepdrix (zepdrix):

So when we're dealing with sine and cosine, \[\large y=asin(\color{royalblue}{b}t)\] Our period, generally labeled as \(\large T\), is given by, \(\large \dfrac{2\pi}{b}\). So what is our \(\large b\) value in this problem? It appears to be the \(\large 4\pi\) yes?

zepdrix (zepdrix):

\[\large T=\frac{2\pi}{4\pi}\]

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

1/2

zepdrix (zepdrix):

Yah sounds right! :) good job.